a.Find the number of three-digit numbers can be formed using the numbers 2,3,5,6,7 and 9.



b. how many are greater than 800




1
Expert's answer
2021-12-15T10:20:42-0500

a. Given that each digit can be used only once, and we want a three-digit number here. Therefore there are 6 ways to fill the first digit (at hundredths place), after this, there are 5 ways to fill the second digit and after filling the 2 digits there would be 1 digit to fill the last one.

Therefore total numbers possible here is computed as:

"= 6 \times 5 \times 4 = 120"

b. The number of numbers greater than 800 is computed here as:

9 _ _

The tens digit here could be filled with 2, 3, 5, 6, or 7 that is 5 ways after which the one digit can be filled with 6 - 2 =4 ways. Therefore total such numbers "= 5 \times 4 = 20"